### v1

@@ C.7 Proof of Theorem 3
{Proof}
In this subsection, we aim to prove that the introduction of the baseline function reduces the sample complexity. From the proof of Theorem 1, the required sample $n_{\min}$ is determined by the upper bound of the sample $\lVert Z_{i}\rVert$. Using (84), we know that this upper bound is probabilistically bounded as follows:
(100) $\begin{split}\lVert\hat{C}_{k}-\hat{b}_{s}(x_{0})\rVert&=\lVert\hat{C}_{k}-b_{s}(x_{0})+b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\\ &\leq\lVert\hat{C}_{k}-b_{s}(x_{0})\rVert+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\end{split}$
Since $\hat{b}_{s}(x_{0})>0$, the term $\lVert\hat{C}_{k}-b_{s}(x_{0})\rVert$ is upper bounded by $\bar{C}-{b}_{s}^{*}(x_{0})$, if ${b}_{s}^{*}(x_{0})\leq\hat{C}_{k}$, and by ${b}_{s}^{*}(x_{0})$, if ${b}_{s}^{*}(x_{0})>\hat{C}_{k}$. Therefore, we have:
(101) $\begin{split}\lVert\hat{C}_{k}-b_{s}(x_{0})\rVert\leq\max\{\hat{C}_{k}-{b}_{s}^{*}(x_{0}),{b}_{s}^{*}(x_{0})\}\leq\max\{\bar{C}-{b}_{s}^{*}(x_{0}),{b}_{s}^{*}(x_{0})\}\end{split}$
@@
(102) $b_{s}(x_{0})\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}(\lVert x_{0}\rVert+(l-1)\lVert\Sigma_{w}\rVert)^{2}.$
If
(103) $\delta_{x}\geq\underline{\delta}_{x}:=\left(1-\frac{n_{x}\lambda_{1}(\Sigma_{w})}{\lVert\Sigma_{w}\rVert^{2}}\right)^{ln_{x}},$
then $b_{s}(x_{0})\leq\bar{C}$, and further $\lVert\hat{C}_{k}-b_{s}(x_{0})\rVert\leq\bar{C}$. From (100), if
(104) $\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\leq\min\{b_{s}(x_{0}),\bar{C}-b_{s}(x_{0})\}=:\epsilon_{v},$
then $\lVert\hat{C}_{k}-b_{s}(x_{0})\rVert\leq\bar{C}$. Therefore, the required sample for variance reduction is at least not larger than the case without variance reduction. Following the proof of Theorem 1, we can obtain the expression of $N_{3}$:
(105) $N_{3}:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{11}(C(K),\delta_{x},\hat{b}_{s}(x_{0}))\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$
@@
and $\alpha_{12}(\hat{b}_{s}(x_{0})):=\frac{n_{x}n_{u}}{r}\{\max\{\hat{C}_{k}-{b}_{s}^{*}(x_{0}),{b}_{s}^{*}(x_{0})\}+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\}$.
@@
(106) $\hat{C}_{k}^{V}\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\sqrt{\frac{n_{x}\lambda_{1}(\Sigma_{w})}{1-(\tilde{\delta}_{x})^{\frac{1}{tn_{x}}}}}\right)^{2}=:\bar{C}^{V}(\tilde{\delta}_{x})$
@@
(107) $\mathbb{E}[\hat{C}_{k}^{V}]\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\lVert\Sigma_{w}\rVert\right)^{2}=:\bar{C}_{E}^{V}.$
@@
(108) $\lVert C^{V}_{k}-\mathbb{E}[\hat{C}_{k}^{V}]\rVert\leq\bar{C}_{E}^{V}+\bar{C}^{V}(\tilde{\delta}_{x})$
@@
(109) $\begin{split}&\lVert\mathbb{E}(C^{V}_{k}C^{V~{}\top}_{k})-{C}_{k}^{V}{C}_{k}^{V~{}\top}\rVert\leq\max_{C^{V}_{k}}(\lVert C^{V}_{k}\rVert_{F})^{2}+\lVert\mathbb{E}[C^{V}_{k}]\rVert^{2}_{F}\leq(\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2}.\end{split}$
Using the matrix concentration inequality [11, Lemma B2.5], we get:
(110) $\tilde{n}_{\min}=\frac{2}{(\epsilon^{\prime}_{v})^{2}}\left((\bar{C}_{E}^{V}+\bar{C}^{V})^{2}+\frac{((\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2})\epsilon_{v}}{3}\right)\log\left[\frac{2}{(\tilde{\delta}_{v})}\right].$
Thus, if the number of rollouts used to estimate the baseline function satisfies $n_{v}\geq\tilde{n}_{\min}(\tilde{\delta}_{x},\delta_{v})$, the sample size required to estimate the gradient with the baseline function is at least not larger than the case without the baseline function, with probability $\delta_{v}=(1-\tilde{\delta}_{v})\tilde{\delta}_{x}$.
C.8 Proof of Theorem 4.3
{Proof}
@@
Then convergence is guaranteed. Observing that: $\hat{K}_{i+1}-K_{i+1}=\eta_{i}(\hat{\nabla}C(\hat{K}_{i})-\nabla C(K_{i}))$, together with Theorem 3.1:
@@
Then we can derive the requirement on the estimation accuracy $\epsilon_{\mathrm{GD}}$ on ^$\hat{\nabla}C(\hat{K}_{i})$
$\epsilon_{\mathrm{GD}}:=\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
Then we should choose the $l,n,r$ as discussed in Theorem 1 to satisfy the accuracy and probability $\delta$ above. The gradient descent algorithm can reach the desired accuracy $\epsilon$ with probability $\delta$. The remainder of the proof follows Theorem 3.1, which assumes that the gradient estimate meets the required accuracy at each iteration, guaranteed by the union bound.
C.9 Proof of Theorem 4.4
{Proof}
@@
^$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
@@
^$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert+\lVert{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert$
From Weyl’s Theorem [10, Theorem 31], if $\epsilon\leq\frac{\lambda_{1}(\Sigma_{w})}{2}$, then it holds that $\lambda_{1}(\hat{\Sigma}_{\hat{K}_{i}})\leq\lambda_{1}({\Sigma}_{w})-\lVert\hat{\Sigma}_{\hat{K}_{i}}-\Sigma_{\hat{K}_{i}}\rVert=\frac{\lambda_{1}({\Sigma}_{w})}{2}$. Because $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{2}{\lambda_{1}(\Sigma_{w})}$, we aim to make sure that
^$\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}=:\epsilon_{NPG}^{C},$
so that ^$\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}$. Then we should choose the $l_{1},n_{1},r_{1}$ in Algorithm 1 according to Theorem 1 to satisfy the accuracy on ^$\hat{\nabla}C(\hat{K}_{i})$ and the probability $\delta^{\frac{1}{2}}$ above.
@@
(111) $\begin{split}\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert&\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\lVert{\nabla}C(\hat{K}_{i})\rVert}\\ &\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}=:\epsilon_{NPG}^{\Sigma}\end{split}$
@@
$\mathbb{P}\left\{\lVert C(\hat{K}_{i+1})-C({K}_{i+1})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\right\}\geq\delta$
Then natural policy gradient can reach the desired accuracy $\epsilon$ with probability $\delta$. The remainder of the proof follows Theorem 19, which assumes that the gradient estimate and covariance matrix meet the required accuracy at each iteration, guaranteed by the union bound.

### v2

@@ C.7 Proof of Theorem 6
Proof 13
In this subsection, we aim to prove that the introduction of the baseline function reduces the sample complexity. From the proof of Theorem 4, the required sample $n_{\min}$ is determined by the upper bound of the sample $\lVert Z_{k}\rVert$. Using (85), we know that this upper bound is probabilistically bounded as follows:
$\begin{split}\lVert\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0})\rVert&=\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})+b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\\ &\leq\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\end{split}$ (101)
Since $\hat{b}_{s}(x_{0})>0$, the term $\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert$ can be bounded as follows. If ${b}_{s}(x_{0})\leq\hat{C}_{\bar{K}_{k}}$, then bounded by $\bar{C}-{b}_{s}(x_{0})$ where $\bar{C}\geq\hat{C}_{\bar{K}_{k}}$ for all $k$ and $\bar{C}$ was introduced in (85). Otherwise, if ${b}_{s}(x_{0})\leq\hat{C}_{\bar{K}_{k}}$, and by ${b}_{s}(x_{0})$, if ${b}_{s}(x_{0})>\hat{C}_{\bar{K}_{k}}$. Therefore, we have:
$\begin{split}\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert\leq\max\{\hat{C}_{\bar{K}_{k}}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}\leq\max\{\bar{C}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}\end{split}$ (102)
@@
$b_{s}(x_{0})\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}(\lVert x_{0}\rVert+(l-1)\lVert\Sigma_{w}\rVert)^{2}.$ (103)
Compare the expression of $b_{s}(x_{0})$ and $\bar{C}$ in (85), we know that $b_{s}(x_{0})\leq\bar{C}$, and further $\lVert\hat{C}_{\bar{K}_{k}}-b_{s}(x_{0})\rVert\leq\bar{C}$. From (101), if
$\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\leq\min\{b_{s}(x_{0}),\bar{C}-b_{s}(x_{0})\}=:\epsilon_{v},$ (104)
then $\lVert\hat{C}_{\bar{K}_{k}}-\hat{b}_{s}(x_{0})\rVert\leq\bar{C}$. Therefore, the required sample for variance reduction is at least not larger than the case without variance reduction. Following the proof of Theorem 4, we can obtain the expression of $N_{3}$:
$N_{3}:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{d}^{2}}\left(\alpha_{7}^{2}(C(K),\delta_{x})+\frac{\alpha_{11}(C(K),\delta_{x},\hat{b}_{s}(x_{0}))\epsilon_{d}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta_{d}}\right],$ (105)
@@
and $\alpha_{12}(\hat{b}_{s}(x_{0})):=\frac{n_{x}n_{u}}{r}\{\max\{\bar{C}-{b}_{s}(x_{0}),{b}_{s}(x_{0})\}+\lVert b_{s}(x_{0})-\hat{b}_{s}(x_{0})\rVert\}$.
@@
$\hat{C}_{k}^{V}\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/l}}\right)^{2}=:\bar{C}^{V}(\tilde{\delta}_{x})$ (106)
@@
$\mathbb{E}[\hat{C}_{k}^{V}]\leq\frac{C(K)}{\lambda_{1}(\Sigma_{w})}\left(L_{0}+(l-1)\lVert\Sigma_{w}\rVert\right)^{2}=:\bar{C}_{E}^{V}.$ (107)
@@
$\lVert C^{V}_{k}-\mathbb{E}[\hat{C}_{k}^{V}]\rVert\leq\bar{C}_{E}^{V}+\bar{C}^{V}(\tilde{\delta}_{x})$ (108)
@@
$\begin{split}&\lVert\mathbb{E}(C^{V}_{k}C^{V~\top}_{k})-{C}_{k}^{V}{C}_{k}^{V~\top}\rVert\leq\max_{C^{V}_{k}}(\lVert C^{V}_{k}\rVert_{F})^{2}+\lVert\mathbb{E}[C^{V}_{k}]\rVert^{2}_{F}\leq(\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2}.\end{split}$ (109)
Using the matrix concentration inequality [10, Lemma B2.5], we get:
$\tilde{n}_{\min}=\frac{2}{(\epsilon^{\prime}_{v})^{2}}\left((\bar{C}_{E}^{V}+\bar{C}^{V})^{2}+\frac{((\bar{C}_{E}^{V})^{2}+(\bar{C}^{V}(\tilde{\delta}_{x}))^{2})\epsilon_{v}}{3}\right)\log\left[\frac{2}{(\tilde{\delta}_{v})}\right].$ (110)
Thus, if the number of rollouts used to estimate the baseline function satisfies $n_{v}\geq\tilde{n}_{\min}(\tilde{\delta}_{x},\delta_{v})$, the sample size required to estimate the gradient with the baseline function is at least not larger than the case without the baseline function, with probability $\delta_{v}=1-(1-\tilde{\delta}_{v})(1-\tilde{\delta}_{x})$.
C.8 Proof of Theorem 7
Proof 14
@@ C.8 Proof of Theorem 7
Then convergence is guaranteed. Observing that: $\hat{K}_{i+1}-K_{i+1}=\eta_{i}(\hat{\nabla}C(\hat{K}_{i})-\nabla C(K_{i}))$, together with Theorem 1:
@@
Then we can derive the requirement on the estimation accuracy $\epsilon_{\mathrm{PGD}}$ on $\hat{\nabla}C(\hat{K}_{i})$
$\epsilon_{\mathrm{PGD}}:=\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
Then we should choose the $l,n,r$ as discussed in Theorem 4 to satisfy the accuracy and probability $\delta$ above. The gradient descent algorithm can reach the desired accuracy $\epsilon$ with probability $1-\delta$. The remainder of the proof follows Theorem 1, which assumes that the gradient estimate meets the required accuracy at each iteration, guaranteed by the union bound and using the fact $(1-\delta)^{n_{\mathrm{PGD}}}\geq 1-\delta n_{\mathrm{PGD}}$.
C.9 Proof of Theorem 8
Proof 15
@@ C.9 Proof of Theorem 8
$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}.$
@@
$\lVert\hat{\nabla}C(\hat{K}_{i})\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\nabla}C(\hat{K}_{i}){\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert+\lVert{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}-{\Sigma}_{\hat{K}_{i}}^{-1}\rVert$
From Weyl’s Theorem [9, Theorem 31], if $\epsilon\leq\frac{\lambda_{1}(\Sigma_{w})}{2}$, then it holds that $\lambda_{1}(\hat{\Sigma}_{\hat{K}_{i}})\leq\lambda_{1}({\Sigma}_{w})-\lVert\hat{\Sigma}_{\hat{K}_{i}}-\Sigma_{\hat{K}_{i}}\rVert=\frac{\lambda_{1}({\Sigma}_{w})}{2}$. Because $\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{2}{\lambda_{1}(\Sigma_{w})}$, we aim to make sure that
$\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{2}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}=:\epsilon_{NPG}^{C},$
so that $\lVert\hat{\nabla}C(\hat{K}_{i})-{\nabla}C(\hat{K}_{i})\rVert\lVert\hat{\Sigma}_{\hat{K}_{i}}^{-1}\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{4h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert}$. Then we should choose the $l_{1},n_{1},r_{1}$ in Algorithm 1 according to Theorem 4 to satisfy the accuracy on $\hat{\nabla}C(\hat{K}_{i})$ and the probability $\delta^{\frac{1}{2}}$ above.
@@
$\begin{split}\lVert\hat{\Sigma}_{\hat{K}_{i}}-{\Sigma}_{\hat{K}_{i}}\rVert&\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\lVert{\nabla}C(\hat{K}_{i})\rVert}\\ &\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}^{3}(\Sigma_{w})}{8h_{\nabla}(C(\hat{K}_{i}))\lVert\Sigma_{K}^{*}\rVert\sqrt{b_{\nabla}(C(\hat{K}_{i}))}}=:\epsilon_{NPG}^{\Sigma}\end{split}$ (111)
@@
$\mathbb{P}\left\{\lVert C(\hat{K}_{i+1})-C({K}_{i+1})\rVert\leq\frac{\epsilon\sigma\lambda_{1}{(R)}\lambda_{1}(\Sigma_{w})}{2\lVert\Sigma_{K}^{*}\rVert}\right\}\geq 1-\delta$
Then natural policy gradient can reach the desired accuracy $\epsilon$ with probability $1-\delta$. The remainder of the proof follows Theorem 2, which assumes that the gradient estimate and covariance matrix meet the required accuracy at each iteration, guaranteed by the union bound and using the fact $(1-\delta)^{n_{\mathrm{NPG}}}\geq 1-\delta n_{\mathrm{NPG}}$.